OpenStudy (anonymous):

Simplify: Sqrt 15x^5 / sqrt 3x

OpenStudy (anonymous):

\[\sqrt{15x ^{5}}\div \sqrt{3x}\]

OpenStudy (anonymous):

\[x^2\sqrt(5)\]

OpenStudy (anonymous):

Don't forget to fan :$